OpenStudy (anonymous):

What are the coordinates of the vertex of y = x2 - 2x - 7?

OpenStudy (campbell_st):

the vertex form of a parabola is \[y = (x - h)^2 + k\] so group your equation as \[y = (x - 2x ) - 7\] complete the square in x

OpenStudy (anonymous):

I have no idea how to do it

OpenStudy (campbell_st):

oops should be \[y = (x^2 -2x ) - 7\]

OpenStudy (campbell_st):

ok... a perfect square is \[(x - a)^2 ~~~ or ~~~~(x + a)^2\] your question is similar the the 1st one. expanding it gives \[(x - a)^2 = x^2 - 2ax + a^2\]

OpenStudy (campbell_st):

so looking at the 1st part of your problem \[y = (x^2 - 2x ) \] what do you think the value of a is.. when you compare

OpenStudy (anonymous):

x^2?

OpenStudy (campbell_st):

look at \[x^2 - 2ax + x^2 ~~~ and ~~~~ x^2 - 2x + ?\] what do you think a is

OpenStudy (anonymous):

7? i dont know

OpenStudy (campbell_st):

here is an alternate method, this is part of the general quadratic formula. for a quadratic \[ax^2 + bx + c\] the line of symmetry is \[x = \frac{-b}{2a}\] this is also the x value in the vertex. so you have a = 1 and b = -2 what is the value of x...?

OpenStudy (jhannybean):

Complete the square!!! mwuahahahaha

OpenStudy (anonymous):

1?

OpenStudy (campbell_st):

well we are at an alternate method that gives the same result fabulous @jwilliams03 you know x = 1 is part of the vertex. to find the y part, substitute x = 1 into the equation. What do you get...?

OpenStudy (anonymous):

1^2 - 2* 1 - 7 right?

OpenStudy (jhannybean):

Vertex form : \(y=(x-h)^2 +k\) Your goal is to change \(y=x^2-2x-7\) to fit that form.\[y=(x^2-2x)-7\]\[y=\left(x^2-2x+1\right) -7+1\]When you simplify that you will get the same thing @campbell_st has gotten :) this is just an alternate way!

OpenStudy (campbell_st):

yes... that's it, what is the value..?

OpenStudy (anonymous):

-8

OpenStudy (campbell_st):

yes @Jhannybean that's how I started by @jwilliams03 stuggled with the concept

OpenStudy (campbell_st):

great so you know x = 1 and y = -8 that's the vertex (1, -8) to easy

OpenStudy (anonymous):

Okay thanks!

OpenStudy (jhannybean):

I caught my mistake, it was meant to read \(y=\left(x^2-2x+1\right) -7-1\)